HUBBLE SPACE TELESCOPE DAILY REPORT #1055 PERIOD COVERED: 0700 EST Daily Status Report as of 018/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed S/C 5254 (GHRS - SMOV Side 1 Activations) The High Resolution Spectrograph (Side One) was turned on (017/0110Z) by ground commanding. No anomalies were noted. The HRS was safed by stored commands (017/0225Z), after a successful Side One low voltage turn-on. As was planned, the HRS was recovered to side two a few minutes later. No problems were seen during the Side One turn-on, safing, or recovery. Subsequently, the Side One DEB was successfully turned on, at three hours intervals. 1.2 Completed FOS/RD/BL/COSTAR 4747 (COSTAR/FOS Fine Alignment - Focus) Both the Faint Object Spectrograph and the COSTAR were used to perform alignment and focus operations on the FOS M1 mirror. Observations and stored command moves (X and Y tilt) were monitored when the telemetry was available. All of the guide star acquisitions and reacquisition were executed as planned. Fine lock was obtained and the peakup operations worked as planned. No anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled :5 Successful :5 2.3 Significant HSTARs: None 2.4 Operations Notes: Nominal operations, no anomalies noted. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete S/C 5254 (GHRS SMOV Side 1 Activation) 3.2 Complete Two Sets of WF/PC-2 4709 (WF/PC-2 Internal Monitor) 3.3 Complete FOC/96 4788 (SMOV COSTAR/FOC Point Spread Function) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: None /DMH -END-